# Nonlinear Fore(Back)casting and Innovation Filtering for Causal-Noncausal VAR Models

### v3

@@ 6.1.2 Estimated Predictive Density, Point and Interval Forecasts
Let us now consider the oos nonlinear forecast one step ahead performed at date $T=590$. At time T=590, the process takes values -3.367 and -0.239. The true values of $Y_{1}$ and $Y_{2}$ at T+1=591 are -2.260 and -0.331, respectively. The nonlinear forecasts are summarized by the predictive density that can be used to compute the pointwise predictions of $Y_{1,T+1}$ and $Y_{2,T+1}$. The predictive bivariate density is estimated and shown in Figure 2 along with the estimated predictive marginal densities of $Y_{1,T+1}$ and $Y_{2,T+1}$, respectively. The predictive density is estimated from formula (3.12) one-step ahead oos by using a kernel estimator over a grid of 100 values below and above $Y_{1}$ and $Y_{2}$, with Gaussian kernels and bandwidths $h_{2}=1$ and $h_{11}=s.d.(\varepsilon_{1}),h_{12}=s.d.(\varepsilon_{2})$ (see On-Line Appendix C for kernel density estimators).
(a) Estimated Joint Predictive Density
(b) $Y_{1,T+1}$
(c) $Y_{2,T+1}$
Figure 2: Estimated Predictive Joint and Marginal Densities
The estimated point forecasts are obtained from the mode of the predictive density. The forecast of $Y_{1,591}$ based on the GCov estimated parameters is -2.80 and the point forecast of $Y_{2,591}$ is -0.30. The estimated prediction intervals at level 0.80 determined from the predictive density are as follows: The estimated prediction interval for $Y_{1,591}$ at level 0.80 is [-4.80, -0.80] and the prediction interval for $Y_{2,591}$ is [-2.60, 2.10]. The rationale for choosing level 80% is to ensure a sufficiently large number of observations in the tails for computing the quantiles of the predictive density. Both prediction intervals contain the true future values of the process.
Next, we compute oos one-step ahead forecasts based on the sub-sample of 500 observations on $Y_{t}$. Figure 3 displays 100 one-step ahead oos forecasts along the trajectory.
(a) Estimated 1-step-ahead Forecasts of $Y_{1,t}$
(b) Estimated 1-step-ahead Forecasts of $Y_{2,t}$
Figure 3: Estimation-Based oos One-Step Ahead Forecasts true:solid line, forecast:dashed line
6.1.3 Unconditional Coverage of the Estimated Prediction Interval
Let us now consider a bivariate VAR(1) with the autoregressive matrix:
$\Phi=\left(\begin{array}[]{cc}0.9&-0.3\\ 0.0&1.2\end{array}\right),$
and eigenvalues $\lambda_{1}=0.9,\lambda_{2}=1.2$, which are closer to the unit circle. The errors $\varepsilon_{t}$ have an identity variance-covariance matrices and t-student distributions with 3, 6 and 9 degrees of freedom.
The simulated bivariate series are of length 100, 500 and 1000 and the DGP is replicated 500 times. The last observation from each simulated path is set aside. It is forecast and used for forecast coverage assessment. The autoregressive parameters $\hat{\Phi}$ are estimated by the GCov estimator from the errors $\varepsilon_{t}=Y_{t}-\Phi_{1}Y_{t-1}$, their second, third and fourth powers, and their lags up to $H=10$. The objective function maximizing algorithm is each time initiated at the starting values 0.1, 0.1, 0.5, 0.5. The oos predictive density given in equation (3.12) of $y(1000)$ is evaluated, given the parameter estimates, over a grid of 100 possible future values for each of the components series. We use Gaussian kernels and bandwidths $h_{2}=s.d.(Y_{2,t})$ $h_{11}=s.d.(\varepsilon_{1})$ and $h_{12}=s.d.(\varepsilon_{2})$ (see On-Line Appendix C for kernel density estimators). The mode of the predictive density estimated from formula (3.12) with Gaussian kernels provides the point forecast. The oos prediction intervals at horizon 1 are obtained from the marginal 10th and 90th percentiles of the bivariate predictive density. As before, the rationale for choosing level 80% is to ensure a sufficiently large number of observations in the tails for computing the quantiles of predictive densities.
The unconditional coverage of the estimated prediction interval is reported in Table 1 below for the sample sizes T=500 and T=1000 and t-student error $\varepsilon_{t}$ distributions with 3, 6 and 9 degrees of freedom.
Table 1: Coverage of GCov Estimated Prediction Interval at 80%
T=100 T=500 T=1000
VAR(1) with eigenvalues 0.9, 1.2
component t(3) t(6) t(9) t(3) t(6) t(9) t(3) t(6) t(9)
$y_{1}(T+1)$ 80.2 82.8 87.4 82.2 86.0 84.2 84.6 84.4 88.6
$y_{2}(T+1)$ 90.4 89.4 90.8 90.6 91.6 92.40 91.4 91.0 92.8
We observe that the coverage is either greater or close to the theoretical size of the prediction interval for all sample sizes. Then, the bootstrap adjustment could be applied.
6.1.4 Estimated Prediction Set Uncertainty
This Section illustrates the uncertainty on the conditional prediction interval described in Section 5.2. We consider the simulated trajectory of length T=200 of the bivariate VAR(1) process with autoregressive matrix, illustrated in Table 1 with eigenvalues $\lambda_{1}=0.9,\lambda_{2}=1.2$ and t(6) distributed errors with an identity variance-covariance matrix.
We are interested in the conditional prediction interval out-of-sample of the first component $Y_{1,T+1}$ = $Y_{1,200}$ given the past and current values of the two series.
The mixed VAR(1) model is estimated by the GCov estimator from the observations $t=1,...,199$ with four power transforms of model errors and lag $H=10$, providing the following estimates of the autoregressive parameters: 0.8931, -0.2146, 0.0180, 1.2797, and the following estimates of eigenvalues: 0.903 and 1.269.
The estimated predictive density of $Y_{200}$ conditional on $Y_{199}$ is computed from formula (3.12) by using kernel smoothed density estimators. More specifically, we employ Gaussian kernels and bandwidths $h_{2}=s.d.(Y_{2}),h_{11}=s.d.(\varepsilon_{1}),h_{12}=s.d(\varepsilon_{2})$ (see On-Line Appendix C for kernel density estimators). For $\alpha_{1}=0.2$ and $\Phi^{-1}(\alpha_{1}/2)=-1.28$, the estimated prediction interval of $Y_{1,T+1}=Y_{1,200}$ is $\widehat{PI}(y,\alpha_{1})=[-6.954,0.360].$ It contains the true value of $Y_{1,200}=-0.696$.
In the next step, we replicate the initial paths $S=50$ times by backcasting from the bivariate terminal condition $Y_{199}=[-1.188,0.473]$. We use the backcasting formula given in Corollary 2, evaluated from estimated model parameters and kernel density estimators. We employ Gausian kernels and bandwidths $h_{1}=s.d.(Y_{1,t}),h_{11}=s.d.(\varepsilon_{1})$ and $h_{12}=s.d.(\varepsilon_{2})$.
The randomness of the backcasted paths is generated as follows: i) We first draw 100 values as if the components $Y_{1,T-1},Y_{2,T-1}$ were independent conditional on $Y_{T}$. This is done by inverting the estimated conditional c.d.f.s of $Y_{1,T-1}$ and $Y_{2,T-1}$ given $Y_{T}$. This provides the sampling with the importance (misspecified) density; ii) Next, we re-sample in the set of values obtained in step i) above, with the weights proportional to the ratio of the joint backward predictive density divided by the product of the two marginal backward predictive densities. This procedure adjusts for the omitted cross-sectional dependence in step i).
The parameter $\Phi$ and functional parameter $g$ are re-estimated from each replicated path, and then the values of $\hat{m}^{s}$, $\hat{\sigma}^{s}$ are computed from the quantiles of $50$ predictive densities of $Y_{T+1}$ conditional on $Y_{T}$. They are next used to compute the bootstrap confidence interval at level $\alpha_{2}=0.1$ of the prediction interval $\widehat{CSPI}(y,\alpha_{1},\alpha_{2})=[-11.839,5.246],$ with the solution $\hat{q}(y,\alpha_{1},\alpha_{2})=2.99$. The confidence interval $\widehat{CSPI}$ of the prediction interval is much larger than the interval $\widehat{PI}$. This result illustrates the importance of taking into account the estimation risk on $\Phi,g$ when providing a prediction interval, especially that the estimator of the functional parameter $g$ converges at a lower speed than the parameter estimator. The effect of estimation risk is twofold: i) the length of the interval has almost doubled, ii) the interval became less symmetric with respect to 0.
6.2 Application to Cryptocurrency Prices
Cryptocurrency rates often display speculative bubbles and can be modelled as mixed VAR models [see e.g. Gourieroux, Hencic (2015) for an early paper on cryptocurrency dynamics and Hall, Jasiak (2024)]. Since 2021, the cryptocurrency market has evolved towards becoming better organized and regulated. First, various cryptocurrency indexes have appeared, including the S&\&&P Bitcoin Index, S&\&&P Ethereum Index and the S&\&&P Cryptocurrency MegaCap Index for tracking separately and jointly the performance of the cryptocurrencies with the highest capitalization, Bitcoin and Ethereum. The current market capitalizations of Bitcoin (BTC) and Ethereum (ETH) are of about 938.3 and 363.41 billion USD, respectively. More recently, several Exchange Traded Funds (ETF) for cryptocurrency became available. On January 10, 2024, the SEC (Securities and Exchange Commission) approved the listing and trading of crypto ETF shares on registered securities exchanges in the US, indirectly opening this market to retail investors  [^13] The OSC (Ontario Securities Commission) approved a Bitcoin ETF in February 2021.. This development marks the beginning of a new episode in the cryptocurrency markets and motivates our empirical study of Bitcoin and Ethereum.
Estimation
Let us consider the bivariate series of Bitcoin and Ethereum prices in US Dollars. The bivariate series of 257 daily adjusted closing BTC/USD and ETH/USD exchange rates recorded between July 21, 2021 and April, 04, 2022  [^14] Data Source: Yahoo Finance Canada https://ca.finance.yahoo.com/ are displayed in Figure 4.
Figure 4: BTC/USD and ETH/USD Daily Closing Rates, July 21, 2021 to April 04, 2022 BTC/USD: solid black line, ETH/USD (times 30): dashed red line
We observe that both series display comovements over time and their dynamics are characterized by spikes and local trends/bubbles.
The bivariate VAR model is estimated from the data rescaled and adjusted by a polynomial function of time of order 2 for the hump-shaped pattern in the middle of the sampling period. This pre-filtering by a deterministic function is used instead of a moving average filter that could create spurious noncausal effects. The estimation is performed without imposing any distributional assumptions on the errors. We use the semi-parametrically efficient GCov estimator with nonlinear transformations including powers two, three and four of the errors summed up to lag 10. The estimated matrix $\hat{\Phi}$ is:
$\hat{\Phi}=\left(\begin{array}[]{cc}-0.0901&1.1998\\ 0.4183&0.7724\end{array}\right),$
with eigenvalues -0.488 and 1.171. Hence, there is a single noncausal component that captures the common bubbles and spikes. The standard errors of $\hat{\Phi}_{1,1}$ and $\hat{\Phi}_{1,2}$ are 0.014 and 0.013. For $\hat{\Phi}_{2,1}$ and $\hat{\Phi}_{2,2}$ the standard errors take values 0.022 and 0.019, respectively. To check the assumption of serial independence of errors $\varepsilon_{t}$, Figures 5 and 6 provide the autocorrelations of the approximated errors and their squares. They indicate that the error is close to a bivariate white noise and the model provides a satisfactory fit.
Figure 5: ACF of $\hat{\varepsilon}_{t}$
Figure 6: ACF of Squared $\hat{\varepsilon}_{t}$
The Jordan decomposition of matrix $\hat{\Phi}$ is: $\hat{\Phi}=\hat{A}\hat{J}\hat{A}^{-1},$ with
$\hat{A}=\left(\begin{array}[]{cc}-3.013&0.951\\ 1&1\end{array}\right)$, $\hat{J}=\left(\begin{array}[]{cc}-0.488&0\\ 0&1.171\end{array}\right)$, $\hat{A}^{-1}=\left(\begin{array}[]{cc}-0.252&0.240\\ 0.252&0.759\end{array}\right)$.
Forecasting
We illustrate the forward prediction method by forecasting the adjusted closing BTC/USD and ETH/USD exchange rates on April 4, 2022  [^15] This last observation is excluded from the estimation. equal to 46622.67 and 3521.24, respectively. The point forecast of demeaned and rescaled $y_{1,T+1}$ is -740.00, and the forecast of demeaned $y_{2,T+1}$ is 105.00. The forecast interval of $y_{1}$ at 0.90% is [-940.00 and -150.00] and contains the true value -844.81. The forecast interval of $y_{2}$ at 0.90% is [57.00, 174.00] and contains the true value 126.68. After adjusting for the mean and scale, we obtain forecasts of 46727.0 for Bitcoin and 3499.56 for Ethereum prices, which are off by 105 and 22 Dollars, respectively, outperforming a combined ”no-change” forecast based on the previous day values.
The state variables
The Jordan decomposition is used to estimate the state variables $Z_{1,t},Z_{2,t}$, which have financial interpretations in the cryptocurrency model. The noncausal state variables are the following combinations of the observed series:
$\hat{Z}_{1,t}=-0.252Y_{1,t}+0.240Y_{2,t},\;\;\hat{Z}_{2,t}=0.252Y_{1,t}+0.759Y_{2,t},$
The noncausal state variable allocates positive weights to Bitcoin and Ethereum, like the MegaCap index. The causal state variable has a negative weight of Bitcoin. It represents a risk-neutral fund, i.e. a fund which is little sensitive to the changes in $Z_{2,t}$, and hedges against the common bubble effects. The two state variables are plotted in Figure 7.
Figure 7: State Variables: black line: causal; red line: noncausal
We observe that the noncausal state variable captures the spikes and explosive patterns in cryptocurrency rates, while the causal state variable is much more stable.
Let us now compare the state variables with the factors obtained from a SVD (Singular Value Decomposition) of the marginal variance-covariance matrix $Var(Y_{t})$. It is equivalent to consider the spectral decomposition of $Var(Y)$ and $Var(Z)$ as $Y$ and $Z$ satisfy a one-to-one linear relationship. It follows from (3.7)-(3.9) applied to the bivariate mixed VAR(1) model that:
$Z_{1,t}=\sum_{h=0}^{\infty}[j_{1}^{h}a^{1}\varepsilon_{t-h}],\;\;\;Z_{2,t}=-\sum_{h=0}^{\infty}[(1/j_{2})^{1+h}a^{2}\varepsilon_{t+h+1}],$
where $a^{1},a^{2}$ are the rows of $A^{-1}$. We deduce that the matrix $Var(Z)$ is diagonal with
$Var(Z_{1,t})=\frac{1}{1-j_{1}^{2}}a^{1}\Sigma(a^{1})^{\prime},\;\;\;Var(Z_{2,t})=\frac{1}{j_{2}^{2}-1}a^{2}\Sigma(a^{2})^{\prime}$
In particular, the ordered eigenvectors are driven by either $Z_{1,t},Z_{2,t}$ (resp. $Z_{2,t},Z_{1,t}$ ) depending if $Var(Z_{1,t})>Var(Z_{2,t})$ (resp. $Var(Z_{1,t})<Var(Z_{2,t})$. The ordering can be reversed depending if the risk due to the explosive patterns is greater than the traditional risk measured by the variance of the causal component, or vice versa. In our example, we get $Var(Z_{2,t})$ = 46904282.5 and $Var(Z_{1,t})$ = 347597.4, as the variation of the explosive component is much higher.
Causal Analysis
i) Conditional Value-at-Risk of the noncausal component
The mixed VAR(1) model has a multiplicity of nonlinear autoregressive representations, and a multiplicity of state variables $Z_{1,t},Z_{2,t}$ because of the multilplicity of Jordan forms. The former multiplicity gets reduced when the shocks on $Z_{1},Z_{2}$ are ordered. This leads to a recursive form of the nonlinear autoregressive model:
$Z_{2,t}$ $=$ $G_{2}(Z_{2,t-1};v_{2,t}),$ (6.1)
$Z_{1,t}$ $=$ $G_{1|2}(Z_{2,t},Z_{1,t-1};v_{1,t})$ (6.2)
where function $G_{2}$ in $Z_{2,t}=G_{2}(Z_{2,t-1};v_{2,t})$ is the inverse of function $v_{2,t}=\Phi^{-1}[F_{2}(Z_{2,t}|Z_{2,t-1})]$ in (4.3). By the Markov property of $(Z_{2,t})$ [see Corollary 4], $Z_{1,t-1}$ does not appear in equation (6.1).
For illustration, we focus on the CBS on the state variable $Z_{2}$, which is a systemic factor of bubble risk, that is on the first equality of the recursive system.
It is easy to see that function $G_{2}(z,v_{2})$ is the conditional Value-at-Risk (VaR) of $Z_{2}$ at level $\alpha=\Phi(v_{2})$. This conditional VaR is easily estimated nonparametrically with a closed form expression of the estimator (see online Appendix D)
ii) Common Bubble Shock (CBS)
Since $\hat{G}(z,v_{2})$ has a simple closed form expression , the equation (6.1) can be used to compare the future paths $Z_{2,T+1}^{s}$, $Z_{2,T+2}^{s},...$ corresponding to a sequence of simulated innovations $(v_{T+1}^{s},...,v_{T+H}^{s})$ and to shocked at T+1 simulated shocks $(v_{T+1}^{s}+\delta,...,v_{T+H}^{s})$.
Let us now consider the effect of a common bubble shock $\delta_{2}$ performed at date $T$ on $z_{2,T}$. Since the dynamic model is nonlinear, the IRF is nonlinear in $\delta_{2}$ and in the value $y=(y_{1},y_{2})^{\prime}$ of $Y_{T}$.
Figure 8: Impulse Response Functions: black line: baseline
We observe that the impulse response functions are not symmetric in $\delta_{2}=-2,-1,0,1,2$

### v4

@@ 6.1 Simulation Study
Let us now consider the oos nonlinear forecast one step ahead performed at date $T=500$ when the process takes values $Y_{1,500}=-3.367$ and $Y_{2,500}=-0.239$. The true values of $Y_{1,501}$ and $Y_{2,501}$ are -2.260 and -0.331, respectively. The nonlinear forecasts are summarized by the predictive density, whose mode provides pointwise predictions of $Y_{1,T+1}$ and $Y_{2,T+1}$. It is estimated from formula (3.2) one-step ahead oos by using a kernel estimator over a grid of 100 values below and above $Y_{1}$ and $Y_{2}$, with Gaussian kernels and bandwidths $h_{2}=1$ and $h_{11}=s.d.(\varepsilon_{1}),h_{12}=s.d.(\varepsilon_{2})$ (see Online Appendix B for kernel density estimators). The estimated point forecasts are $\hat{Y}_{1,501}=-2.80$ and $\hat{Y}_{2,501}=-0.30$. The estimated prediction intervals at level 0.80 determined from the predictive density are [-4.80, -0.80] for $Y_{1,501}$ and [-2.60, 2.10] for $Y_{2,501}$. The rationale for choosing level 80% is to ensure a sufficiently large number of observations in the tails for reliable estimation of the quantiles of predictive density. Both prediction intervals contain the true future values of the process. Additional results on the coverage of the estimated prediction interval and on the estimated prediction set uncertainty are provided in Online Appendix D.
Figure 2: State variables: $\hat{Z}_{1}$: solid line, $\hat{Z}_{2}$: dashed line
6.2 Application to Real Oil Prices and Real GDP Growth Rates
In this Section, we apply the mixed VAR(1) model to analyse jointly the real oil prices and the real GDP growth rates.
a) The data
We examine the quarterly series of oil prices and US GDP growth rates over the period: Q1 1986 - Q2 2019. The real US GDP growth rate series is calculated from Real Gross Domestic Product, Quarterly, Seasonally Adjusted Annual Rate available at at https://fred.stlouisfed.org from the Federal Reserve Economic Data.
The oil prices are provided online by the US Energy Information Administration under the Short-Term Energy Outlook Real and Nominal Prices, March 2023 111111 called Quarterly Average Imported Crude Oil Price/barrel, Real Price, deflated by the US Consumer Price CPI index.. This series approximates ”the price of oil paid by US refiners for crude oil purchased from abroad” examined by Kilian and Vigfusson (2017). The series of GDP rates and oil prices (divided by 10) of length T=134 are displayed in Figure 3. We observe, for example, in year 2008, a pronounced bubble in oil prices with a strong negative impact on the GDP growth rate.
b) Estimation
We estimate the causal-noncausal VAR(1) from the demeaned series of GDP growth rates (series 1) and demeaned oil prices divided by 10 (series 2). The sample mean of growth rates is 0.640 and the mean of rescaled oil prices is 6.348. The estimated autoregressive matrix is $\hat{\Phi}=\left[\begin{array}[]{cc}0.2987&-0.0229\\ -0.1527&1.0690\end{array}\right]$ with standard deviations of autoregressive coefficients of 0.014, 0.001, 0.032, 0.004, respectively. The eigenvalues are 0.294 and 1.073. The densities of estimated errors provided in Online Appendix E, Figure a.5 are non-Gaussian.
The presence of a noncausal root reflects the nonlinear dynamic features such as the spikes and a bubble observed in Figure 3. The spikes and bubble in oil prices are accommodated by the strictly stationary mixed VAR(1) model of GDP rates and oil price levels.
c) Prediction
We perform out-of-sample (oos) predictions from the mixed VAR(1) at the points indicated in Figure 3 below, which mark the bubble episode at the end of the sample to ensure a sufficiently large number of prior observations for estimation.
Figure 3: Quarterly Real GDP Rate (black) and Oil Price (red), Q1 1986 -Q2 2019
During the selected period, both series displayed several small sudden changes, before the oil price dropped. The predictive densities are evaluated at each point over a grid of 200 points, equidistant by 0.1 below and above the last observed value of each variable. We use again Gaussian kernels and bandwidths $h_{2}=s.d.(Z_{2})$, $h_{11}=s.d.(\varepsilon_{1})$, and $h_{12}=s.d.(\varepsilon_{2})$ to estimate one-step ahead oos predictive density (3.2) [See Online Appendix E, Figures a.7-a.11 for predictive density plots which vary in $T$121212because of the increasing information set and the dependence on the current environment of $Y_{T}$. and become bi-modal on-bubble when the probability of crash increases].
T+1 forecast interval growth interval oil true values
growth oil q11 q12 q21 q22 growth oil
111 -0.037 6.937 -1.437 0.962 4.337 8.537 -0.501 6.335
113 -0.015 7.333 -1.315 0.784 4.933 8.833 0.070 5.649
115 -0.359 6.221 -1.559 0.540 3.921 7.921 0.643 6.237
116 0.145 6.451 -1.054 1.045 4.051 8.251 0.524 5.595
117 0.121 5.862 -1.078 1.021 3.562 7.562 -0.191 2.765
119 0.077 -0.379 -0.922 1.077 -2.479 1.420 -0.060 0.803
Table 1: On-bubble Prediction Intervals at 80% conditioned on data at $T=110:118$
The predictions (col.2) and prediction intervals (col.3:4) computed oos one-step ahead at selected dates on-bubble for the future values at times T+1 given in col.1
Table 1 reports the point forecasts obtained from the main modes of predictive densities, and the forecast intervals obtained from the sample quantiles of estimated predictive densities. The forecast interval is at level 80% to ensure a sufficient number of observations to estimate the quantiles of the predictive density and to alleviate the potential effect of bimodality. We also consider the predictions of the last 2 points in the trajectory. The results in Table 2 are computed according to the procedure described earlier, starting from the conditioning time T=132 when growth rate $Y_{1,132}=-0.458$ and oil $Y_{2,132}=0.268$ are closer to their averages.
T+1 forecast interval growth interval oil true values
growth oil q11 q12 q21 q22 growth oil
133 -0.099 0.340 -1.099 0.800 -1.759 2.140 -0.097 0.257
134 -0.020 0.307 -1.020 0.879 -1.792 2.007 0.032 1.103
Table 2: Off-bubble Prediction Intervals at 80%, $T=132:133$
The predictions (col.2) and prediction intervals (col.3:4) computed oos one-step ahead off-bubble for the future values at times T+1 given in col.1
These prediction intervals are conditional and depend on the date and distance in time from a bubble. We observe that the prediction intervals on-bubble are longer than off-bubble. In addition, we use the approach outlined in Section 4 to estimate the confidence set of the prediction interval for $Y_{1,134}$. It is based on 100 backcast replications of the series conditional on $Y_{T}$. The confidence set of prediction interval for the demeaned growth rate at level $0.951-\alpha_{2}=$ 0.95 is: $\widehat{CSPI}(y,\alpha_{1},\alpha_{2})=[-1.546,1.405]$. This interval is at a level higher than the prediction intervals in Table 2, accounts also for the estimation risk and is longer.
d) Bubble shock
Let us now examine the effect of a set of the following values of the CBS: $\delta_{2}=-2,-1,0,1,2$ to $z_{2,T}$ performed at date $T=110$ during the bubble episode and at time $T=133$ after the bubble, under the identifying assumption of a recursive form (5.3)-(5.4). Since $Z_{t}=A^{-1}Y_{t}$, that shock can originate from either the oil prices, or growth rates, or both.
(a) on bubble T=110
(b) off bubble, T=133
Figure 4: Shocks to $Z_{2}$, black line: baseline $\delta=0$, colored: response functions: $\delta=+1$: red (thin dashed), $\delta=+2$: bright blue (thick dashed), $\delta=-1$: green (dotted) $\delta=-2$ dark blue: (dash-dotted)
The impulse responses in Figure 4 are computed conditionally on the path of the process up to and including date $T$. Because the dynamic model is nonlinear, the IRF is nonlinear in $\delta_{2}$, and the shock effects need to be compared with a baseline. The baseline path of $Z_{2,T+1}^{b},...,Z_{2,T+10}^{b}$ is computed from 10 random values of standard Normal used as the future $v_{2,T+1},...,v_{2,T+10}$. Next, a shock $\delta$ is added to $v_{2,T+1}$ and the consecutive values of shocked $Z_{2,T+1}^{\delta},...,Z_{2,T+10}^{\delta}$ are calculated recursively by inverting the conditional c.d.f. [See Online Appendix C].
To interpret the shocks, we compare the relative size of shock effects to the baseline, and we conclude that a shock has dissipated when the shocked path overlaps with the baseline. We observe that, on-bubble, shocks to $Z_{2}$ are more long-lasting and more symmetric around the baseline. Off-bubble, we observe that a shock of $\delta=-2$ dissipates much slower than other shocks. Next, we perform shocks of the same size to the causal component $Z_{1}$ using a similar approach (Figure 5). We find that shocks to $Z_{1}$ have weaker effects and dissipate quickly, which is consistent with the fact that the bubble is driven mainly by the noncausal component $Z_{2}$.
(a) on bubble, T=110
(b) off bubble, T=133
Figure 5: Shocks to $Z_{1}$, black line: baseline $\delta=0$, colored: response functions: $\delta=+1$: red (thin dashed), $\delta=+2$: bright blue (thick dashed), $\delta=-1$: green (dotted) $\delta=-2$ dark blue: (dash-dotted)
The shocked component $Z_{2}$ can be combined with the values of causal component $Z_{1}$ conditional on its own past and the shocked values of $Z_{2}$. This is done by predicting $Z_{1}$ from the mode of the conditional density $l(z_{1,t}|z_{2,t},z_{1,t-1})$ given in (5.8) and kernel estimated. Then, we can approximate the shocks to $y_{1}$ and $y_{2}$ as $Y^{\delta}=AZ^{\delta}$. For illustration, Figure 6 displays the shocked values of $Y_{2,T+1}^{b},...,Y_{2,T+10}^{b}$.
(a) on bubble, T=110
(b) off bubble, T=133
Figure 6: Shocks to $y_{2}$, black line: baseline $\delta=0$, colored: response functions: $\delta=+1$: red (thin dashed), $\delta=+2$: bright blue (thick dashed), $\delta=-1$: green (dotted) $\delta=-2$ dark blue: (dash-dotted)
We find that $Y_{2}$ is more responsive to on-bubble shocks. Off-bubble, the shocks dissipate faster except for the large positive shock, which again takes longer to disappear.